OpenStudy (anonymous):

The problem is a^2-2ab-3b^2, the chapter is on Polynomials and Factoring, Here is my work. Can you please tell me if it is right, or where I might of went wrong? a^2-2ba-3b^2 3b-1b=-2b (a-1b)(a+) F=a^2 O=3a I= -1ba L= -3b a^2+3a-1ba-3b

myininaya (myininaya):

what are you trying to do? factor? Ax^2+Bx+C A=1 B=-2b C=-3b^2 Find two factors of A*C that add up to be B A*C=1(-3b^2)=-3b^2=-3b(b) B=-2b=-3b+b so we replace -2ab with -3ab+ab and we have a^2-3ab+ab-3b^2 now factor by grouping a(a-3b)+b(a-3b) (a-3b)(a+b)

OpenStudy (anonymous):

The directions are: Factor completely Remember to look first for a common factor. Check my multiplying. If a polynomial is prime, state this.

myininaya (myininaya):

any questions?

OpenStudy (anonymous):

Yeah, why is mine the wrong way? And to me it is all confusing, where did you get-2ab?

OpenStudy (anonymous):

why not b and c?

myininaya (myininaya):

what?

myininaya (myininaya):

we have Aa^2+Ba+C what is A= what is B= what is C= now we want to find two factors of A*C that multiplied to be A*C and add up to be B.